IF A+B=10 AND A^2+B^2=58 THEN WHAT WILL BE VALUE OF A^3+B^3+C^3

Grade:9

1 Answers

Aditya sashidhar
6 years ago
As you said a+b =10 then let we think a=7 b=3 then A^2+ B^2= 58so the answer I think is 27+343+c^3=370+c^3